OpenStudy (anonymous):

A bag has 2 yellow 7 red and 1 green marble what is the probability of picking a green marble replacing it and then picking a yellow marble? 1.1/25 2.1/50 3.49/100 4.7/100

OpenStudy (anonymous):

P(green) = green/total = 1/10 P(yellow) = 2/10 P(green,yellow) = (1/10)(2/10) = (2/100) =(1/50)

OpenStudy (anonymous):

thanks

OpenStudy (mathmale):

This is sampling with replacement. Assuming that the marbles are mixed frequently, one sampling session is independent of any other. So this problem boils down to determining the probabilities of (1) picking a green marble and (2) picking a yellow marble. As Rae DS has already shown, we need only multiply these two probabilities together to obtain the desired joint probability.

OpenStudy (dan815):



OpenStudy (dan815):